# Sequence Reconstruction for Single-Deletion Single-Substitution Channel

### v1

@@ III-B For $\Lambda^{L}_{1}(\bm{x},\bm{x}^{\prime})$ and $\Lambda^{R}_{1}(\bm{x},\bm{x}^{\prime})$
For each $i\in\{1,2,3\}$, let $\Lambda^{L}_{1,i}=\Lambda^{L}_{1,i}(\bm{x},\bm{x}^{\prime})$ be the set of all $(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})\in\Lambda^{L}_{1}(\bm{x},\bm{x}^{\prime})$, where $(j,j^{\prime})\in[n]\times[n]$ and $j\leq j^{\prime}$, such that the conditions of Case $i$ hold. Clearly,
$\Lambda^{L}_{1}=\bigcup_{i=1}^{3}\Lambda^{L}_{1,i}.$
@@
Claim 1.1: If $T^{L}\cap[i_{2}+1,i_{d}]\neq\emptyset$, then $\Lambda^{L}_{1,1}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $T^{L}\cap[i_{2}+1,i_{d}]=\emptyset$, then $\Lambda^{L}_{1,1}=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash i_{d}})\}$.
Claim 1.2: If $|T^{L}\cap[i_{1}+1,i_{d}]|\geq 2$, then $\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $|T^{L}\cap[i_{1}+1,i_{d}]|=1$, then $\Lambda^{L}_{1,2}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})\}$; if $|T^{L}\cap[i_{1}+1,i_{d}]|=0$, then we have $\Lambda^{L}_{1,2}\subseteq\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\}$. [^1] Here $\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})\subseteq\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\}$ means that if $k_{1}~{}($resp. $k^{\prime}_{1})$ exists, then $(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}})\in\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})~{}\big{(}$resp. $(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\in\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})\big{)}$. The usage of the notation $\subseteq$ in Claims 1.2, 2.2, 2.4, 2.6, 2.2, 2.4 and 2.6 should be understood similarly.
Claim 1.3: If $T^{L}\cap[i_{1}+1,i_{d-1}]\neq\emptyset$, then $\Lambda^{L}_{1,3}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $T^{L}\cap[i_{1}+1,i_{d-1}]=\emptyset$, then $\Lambda^{L}_{1,3}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d-1}})\}$.
Similarly, we can divide $\Lambda^{R}_{1}=\Lambda^{R}_{1}(\bm{x},\bm{x}^{\prime})$ into three subsets $\Lambda^{R}_{1,i}=\Lambda^{R}_{1,i}(\bm{x},\bm{x}^{\prime})$, $i=1,2,3~{}($see Table 1), and we can obtain
$\Lambda^{R}_{1}=\bigcup_{i=1}^{3}\Lambda^{R}_{1,i}.$
If $T^{R}\cap[1,i_{1}]\neq\emptyset$, let
@@
Claim 1.1: If $T^{R}\cap[i_{2}+1,i_{d}]\neq\emptyset$, then $\Lambda^{R}_{1,1}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $T^{R}\cap[i_{2}+1,i_{d}]=\emptyset$, then $\Lambda^{R}_{1,1}=\{(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{2}})\}$.
Claim 1.2: If $|T^{R}\cap[i_{1}+1,i_{d}]|\geq 2$, then $\Lambda^{R}_{1,2}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $|T^{R}\cap[i_{1}+1,i_{d}]|=1$, then $\Lambda^{R}_{1,2}=\{(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{1}})\}$; if $|T^{R}\cap[i_{1}+1,i_{d}]|=0$, then we have $\Lambda^{R}_{1,2}(\bm{x},\bm{x}^{\prime})\subseteq\{(x_{[n]\backslash i_{d},x^{\prime}_{[n]\backslash m_{1}-1}}),(x_{[n]\backslash m^{\prime}_{1}},x^{\prime}_{[n]\backslash i_{1}})\}$.
Claim 1.3: If $T^{R}\cap[i_{1}+1,i_{d-1}]\neq\emptyset$, then $\Lambda^{R}_{1,3}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $T^{R}\cap[i_{1}+1,i_{d-1}]=\emptyset$, then $\Lambda^{R}_{1,3}=\{(x_{[n]\backslash i_{d-1}},x^{\prime}_{[n]\backslash i_{1}})\}$.
@@
Table 1. For each $X\in\{L,R\}$, the set $\Lambda^{X}_{1}$ is divided into three subsets $\Lambda^{X}_{1,i}$, $i=1,2,3$, and the set $\Lambda^{X}_{2}$ is divided into six subsets $\Lambda^{X}_{2,i}$, $i=1,2,\cdots,6$, according to the value of $(|S\cap[1,j-1]|,|T^{X}\cap[j+1,j^{\prime}]|,|S\cap[j^{\prime}+1,n]|)$.
@@
Then
$\Omega^{X}_{1}=\bigcup_{i=1}^{3}\Omega^{X}_{1,i}.$
Moreover, we can easily obtain $\Omega^{X}_{1,i}$ from $\Lambda^{X}_{1,i}$ by Remark 3. As an example, consider $\Omega^{R}_{1,2}$ with $|T^{R}\cap[i_{1}+1,i_{d}]|=1$. By Claim 1.2, we have $\Lambda^{R}_{1,2}=\{(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{1}})\}$. Let $\{j_{1}^{\prime}\}=T^{R}\cap[i_{1}+1,i_{d}]$. Then by Remark 3, we can obtain $\Omega^{R}_{1,2}=B^{\text{S}}_{1}(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{1}})=\{\phi^{i_{d}}_{j_{1}^{\prime}-1;a}(\bm{x}):a\in\Sigma_{q}\}=\{\phi^{i_{1}}_{j_{1}^{\prime};a}(\bm{x}^{\prime}):a\in\Sigma_{q}\}$. In particular, $|B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})|\leq q$ for each $(\bm{z},\bm{z}^{\prime})\in\Lambda^{X}_{1}$ and each $X\in\{L,R\}$.
@@
Let $k_{a}=\max(T^{L}\cap[1,i_{1}]\backslash k_{1})$ if $|T^{L}\cap[1,i_{1}]|\geq 2$, and $k_{a}=1$ otherwise; let $k^{\prime}_{a}=\min(T^{L}\cap[i_{d}+1,n]\backslash\{k^{\prime}_{1}\})-1$ if $|T^{L}\cap[i_{d}+1,n]|\geq 2$, and $k^{\prime}_{a}=n$ otherwise. Then we have
$k_{a}<k_{1}\leq i_{1}<i_{d}<k^{\prime}_{1}\leq k^{\prime}_{a}.$ (11)
@@
Suppose $(j,j^{\prime})$ satisfying the conditions of Case 1. Then from $|S\cap[1,j-1]|=1$ and $|S\cap[j^{\prime}+1,n]|=0$, we have
$i_{1}<j\leq i_{2}\leq i_{d}\leq j^{\prime}.$
Combining this with $|T^{L}\cap[j+1,j^{\prime}]|=0$ (condition of Case 1), we obtain $T^{L}\cap[i_{2}+1,i_{d}]\subseteq T^{L}\cap[j+1,j^{\prime}]=\emptyset$. Hence, if $T^{L}\cap[i_{2}+1,i_{d}]\neq\emptyset$, then we have $\Lambda^{L}_{1,1}(\bm{x},\bm{x}^{\prime})=\emptyset$.
Conversely, suppose $T^{L}\cap[i_{2}+1,i_{d}]=\emptyset$. We need to prove that $\Lambda^{L}_{1,1}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash i_{d}})\}$. Let $j_{1}=\max(T^{L}\cap[i_{1}+1,i_{2}])$ if $T^{L}\cap[i_{1}+1,i_{2}]\neq\emptyset$, and $j_{1}=i_{1}+1$ otherwise. Then by (11), we have
$i_{1}<j_{1}\leq i_{2}\leq i_{d}<k^{\prime}_{1}.$
It is not hard to verify that $(j,j^{\prime})$ satisfies the conditions of Case 1 ((((i.e., $|S\cap[1,j-1]|=1$ and $|T^{L}\cap[j+1,j^{\prime}]|=|S\cap[j^{\prime}+1,n]|=0)$ if and only if
@@
Suppose $(j,j^{\prime})$ satisfying the conditions of Case 2. By the condition $|S\cap[1,j-1]|=|S\cap[j^{\prime}+1,n]|=0$, we have
$j\leq i_{1}<i_{d}\leq j^{\prime}.$
Combining this with $|T^{L}\cap[j+1,j^{\prime}]|=1$ (condition of Case 2), we have $|T^{L}\cap[i_{1}+1,i_{d}]|\leq|T^{L}\cap[j+1,j^{\prime}]|=1$. Hence, $\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})=\emptyset$ if $|T^{L}\cap[i_{1}+1,i_{d}]|\geq 2$.
Conversely, suppose $|T^{L}\cap[i_{1}+1,i_{d}]|\leq 1$. We have the following two cases (Cases a--b).
Case a: $|T^{L}\cap[i_{1}+1,i_{d}]|=1$. Then by the conditions of Case 2, and by (11), we have $k_{1}\leq j\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}$. Similar to Claim 1.1, we can prove $x_{[k_{1},i_{1}]}~{}($resp. $x^{\prime}_{[i_{d},k^{\prime}_{1}-1]})$ is contained in a run of $\bm{x}~{}($resp. $\bm{x}^{\prime})$, so we have $\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{1}\leq j\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}\}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})\}$.
Case b: $|T^{L}\cap[i_{1}+1,i_{d}]|=0$. In this case, by conditions of Case 2, and by (11), we have two possibilities: i) $k_{a}\leq j<k_{1}\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}$, which implies $(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}})\in\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})$; and ii) $k_{1}\leq j\leq i_{1}<i_{d}<k^{\prime}_{1}\leq j^{\prime}\leq k^{\prime}_{a}$, which implies $(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\in\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})$. Thus, we have $\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})\subseteq\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\}$.
@@
Clearly, $(j,j^{\prime})$ satisfies the conditions of Case 3 ((((i.e., $|S\cap[j^{\prime}+1,n]|=1$ and $|S\cap[1,j-1]|=|T^{L}\cap[j+1,j^{\prime}]|=0)$ if and only if

### v2

@@ III-B For $\Lambda^{L}_{1}(\bm{x},\bm{x}^{\prime})$ and $\Lambda^{R}_{1}(\bm{x},\bm{x}^{\prime})$
For each $i\in\{1,2,3\}$, let $\Lambda^{L}_{1,i}=\Lambda^{L}_{1,i}(\bm{x},\bm{x}^{\prime})$ be the set of all $(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}})\in\Lambda^{L}_{1}(\bm{x},\bm{x}^{\prime})$, where $(j,j^{\prime})\in[n]\times[n]$ and $j\leq j^{\prime}$, such that the conditions of Case $i$ hold. Clearly, we have $\Lambda^{L}_{1}=\bigcup_{i=1}^{3}\Lambda^{L}_{1,i}.$
@@
Claim 1.1: If $T^{L}\cap[i_{2}+1,i_{d}]\neq\emptyset$, then $\Lambda^{L}_{1,1}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $T^{L}\cap[i_{2}+1,i_{d}]=\emptyset$, then $\Lambda^{L}_{1,1}=\big{\{}(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash i_{d}})\big{\}}$.
Claim 1.2: If $|T^{L}\cap[i_{1}+1,i_{d}]|\geq 2$, then $\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $|T^{L}\cap[i_{1}+1,i_{d}]|=1$, then $\Lambda^{L}_{1,2}=\big{\{}(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})\big{\}}$; if $|T^{L}\cap[i_{1}+1,i_{d}]|=0$, then we have $\Lambda^{L}_{1,2}\subseteq\big{\{}(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\big{\}}$. [^1] Here $\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})\subseteq\big{\{}(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\big{\}}$ means that if $k_{1}~{}($resp. $k^{\prime}_{1})$ exists, then $(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}})\in\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})~{}\big{(}$resp. $(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\in\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})\big{)}$. The usage of the notation $\subseteq$ in Claims 1.2, 2.2, 2.4, 2.6, 2.2, 2.4 and 2.6 should be understood similarly.
Claim 1.3: If $T^{L}\cap[i_{1}+1,i_{d-1}]\neq\emptyset$, then $\Lambda^{L}_{1,3}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $T^{L}\cap[i_{1}+1,i_{d-1}]=\emptyset$, then $\Lambda^{L}_{1,3}=\big{\{}(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d-1}})\big{\}}$.
Similarly, we can divide $\Lambda^{R}_{1}=\Lambda^{R}_{1}(\bm{x},\bm{x}^{\prime})$ into three subsets $\Lambda^{R}_{1,i}=\Lambda^{R}_{1,i}(\bm{x},\bm{x}^{\prime})$, $i=1,2,3$, according to the value of $(|S\cap[1,j-1]|,|T^{R}\cap[j+1,j^{\prime}]|,|S\cap[j^{\prime}+1,n]|)~{}($see Table 1), and we can obtain $\Lambda^{R}_{1}=\bigcup_{i=1}^{3}\Lambda^{R}_{1,i}.$ If $T^{R}\cap[1,i_{1}]\neq\emptyset$, let
@@
Claim 1.1: If $T^{R}\cap[i_{2}+1,i_{d}]\neq\emptyset$, then $\Lambda^{R}_{1,1}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $T^{R}\cap[i_{2}+1,i_{d}]=\emptyset$, then $\Lambda^{R}_{1,1}=\big{\{}(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{2}})\big{\}}$.
Claim 1.2: If $|T^{R}\cap[i_{1}+1,i_{d}]|\geq 2$, then $\Lambda^{R}_{1,2}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $|T^{R}\cap[i_{1}+1,i_{d}]|=1$, then $\Lambda^{R}_{1,2}=\big{\{}(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{1}})\big{\}}$; if $|T^{R}\cap[i_{1}+1,i_{d}]|=0$, then we have $\Lambda^{R}_{1,2}(\bm{x},\bm{x}^{\prime})\subseteq\big{\{}(x_{[n]\backslash i_{d},x^{\prime}_{[n]\backslash m_{1}-1}}),(x_{[n]\backslash m^{\prime}_{1}},x^{\prime}_{[n]\backslash i_{1}})\big{\}}$.
Claim 1.3: If $T^{R}\cap[i_{1}+1,i_{d-1}]\neq\emptyset$, then $\Lambda^{R}_{1,3}(\bm{x},\bm{x}^{\prime})=\emptyset$; if $T^{R}\cap[i_{1}+1,i_{d-1}]=\emptyset$, then $\Lambda^{R}_{1,3}=\big{\{}(x_{[n]\backslash i_{d-1}},x^{\prime}_{[n]\backslash i_{1}})\big{\}}$.
@@
Table 1. For each $X\in\{L,R\}$ and each $\ell\in\{0,1,2\}$, by Definition 1 and Observation 2, the set $\Lambda^{X}_{\ell}$ can be determined by the tuple $(|S\cap[1,j-1]|,|T^{X}\cap[j+1,j^{\prime}]|,|S\cap[j^{\prime}+1,n]|)$ for each $(j,j^{\prime})\in[n]\times[n]$ such that $j\leq j^{\prime}$. Moreover, the set $\Lambda^{X}_{1}$ is divided into three subsets $\Lambda^{X}_{1,i}$, $i=1,2,3$, and the set $\Lambda^{X}_{2}$ is divided into six subsets $\Lambda^{X}_{2,i}$, $i=1,2,\cdots,6$, according to the value of $(|S\cap[1,j-1]|,|T^{X}\cap[j+1,j^{\prime}]|,|S\cap[j^{\prime}+1,n]|)$.
@@
Then we have $\Omega^{X}_{1}=\bigcup_{i=1}^{3}\Omega^{X}_{1,i}.$ Moreover, we can easily obtain $\Omega^{X}_{1,i}$ from $\Lambda^{X}_{1,i}$ by Remark 3. As an example, consider $\Omega^{R}_{1,2}$ with $|T^{R}\cap[i_{1}+1,i_{d}]|=1$. By Claim 1.2, we have $\Lambda^{R}_{1,2}=\{(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{1}})\}$. Let $\{j_{1}^{\prime}\}=T^{R}\cap[i_{1}+1,i_{d}]$. Then by Remark 3, we can obtain $\Omega^{R}_{1,2}=B^{\text{S}}_{1}(x_{[n]\backslash i_{d}},x^{\prime}_{[n]\backslash i_{1}})=\{\phi^{i_{d}}_{j_{1}^{\prime}-1;a}(\bm{x}):a\in\Sigma_{q}\}=\{\phi^{i_{1}}_{j_{1}^{\prime};a}(\bm{x}^{\prime}):a\in\Sigma_{q}\}$. In particular, we have $|B^{\text{S}}_{1}(\bm{z},\bm{z}^{\prime})|=q$ for each $(\bm{z},\bm{z}^{\prime})\in\Lambda^{X}_{1}$ and each $X\in\{L,R\}$.
@@
Let
$k_{b}=\!\left\{\!\begin{aligned} &\max(T^{L}\cap[1,i_{1}]\backslash k_{1}),~{}\text{if}~{}|T^{L}\cap[1,i_{1}]|\geq 2;\\ &1,~{}\text{otherwise}.\end{aligned}\right.$ (11)
and
$k^{\prime}_{b}=\!\left\{\!\begin{aligned} &\min(T^{L}\cap[i_{d}+1,n]\backslash\{k^{\prime}_{1}\})-1,~{}\text{if}~{}|T^{L}\cap[i_{d}+1,n]|\geq 2;\\ &n,~{}\text{otherwise}.\end{aligned}\right.$ (12)
Then we have
$k_{b}<k_{1}\leq i_{1}<i_{d}<k^{\prime}_{1}\leq k^{\prime}_{b}.$ (13)
@@
Suppose $(j,j^{\prime})$ satisfying the conditions of Case 1. Then from $|S\cap[1,j-1]|=1$ and $|S\cap[j^{\prime}+1,n]|=0$, we have $i_{1}<j\leq i_{2}\leq i_{d}\leq j^{\prime}.$ Combining this with $|T^{L}\cap[j+1,j^{\prime}]|=0$ (condition of Case 1), we obtain $T^{L}\cap[i_{2}+1,i_{d}]\subseteq T^{L}\cap[j+1,j^{\prime}]=\emptyset$. Hence, if $T^{L}\cap[i_{2}+1,i_{d}]\neq\emptyset$, then we have $\Lambda^{L}_{1,1}(\bm{x},\bm{x}^{\prime})=\emptyset$.
Conversely, suppose $T^{L}\cap[i_{2}+1,i_{d}]=\emptyset$. We need to prove that $\Lambda^{L}_{1,1}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash i_{2}},x^{\prime}_{[n]\backslash i_{d}})\}$. Let $j_{1}=\max(T^{L}\cap[i_{1}+1,i_{2}])$ if $T^{L}\cap[i_{1}+1,i_{2}]\neq\emptyset$, and $j_{1}=i_{1}+1$ otherwise. Then by (13), we have $i_{1}<j_{1}\leq i_{2}\leq i_{d}<k^{\prime}_{1}.$ It is not hard to verify that $(j,j^{\prime})$ satisfies the conditions of Case 1 ((((i.e., $|S\cap[1,j-1]|=1$ and $|T^{L}\cap[j+1,j^{\prime}]|=|S\cap[j^{\prime}+1,n]|=0)$ if and only if (see Fig. 3)
@@
Figure 3: An illustration of the pair $(j,j^{\prime})$ in the proof of Claim 1.1. Here, $S=\{i_{1},i_{2},i_{3},i_{4}\}$, $k_{1}$ and $k_{1}^{\prime}$ are defined by (7) and (8) respectively. According to the proof of Claim 1.1, $j_{1}=\max(T^{L}\cap[i_{1}+1,i_{2}])$. We can see that $(j,j^{\prime})$ satisfies the conditions $|S\cap[1,j-1]|=1$ and $|T^{L}\cap[j+1,j^{\prime}]|=|S\cap[j^{\prime}+1,n]|=0$ if and only if it satisfies (14), that is, $j_{1}\leq j\leq i_{2}\leq i_{d}\leq j^{\prime}<k^{\prime}_{1}$. In fact, we have $S\cap[1,j-1]=S\cap[1,j-1]=\{i_{1}\}$ and $T^{L}\cap[j+1,j^{\prime}]=S\cap[j^{\prime}+1,n]=\emptyset$. Moreover, we can see that $x_{[j_{1},i_{2}]}$ is contained in a run of $\bm{x}$ and $x^{\prime}_{[i_{d},k^{\prime}_{1}-1]}$ is contained in a run of $\bm{x}^{\prime}$.
@@
Suppose $(j,j^{\prime})$ satisfying the conditions of Case 2. By the condition $|S\cap[1,j-1]|=|S\cap[j^{\prime}+1,n]|=0$, we have $j\leq i_{1}<i_{d}\leq j^{\prime}.$ Combining this with $|T^{L}\cap[j+1,j^{\prime}]|=1$ (condition of Case 2), we have $|T^{L}\cap[i_{1}+1,i_{d}]|\leq|T^{L}\cap[j+1,j^{\prime}]|=1$. Hence, $\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})=\emptyset$ if $|T^{L}\cap[i_{1}+1,i_{d}]|\geq 2$.
Conversely, suppose $|T^{L}\cap[i_{1}+1,i_{d}]|\leq 1$. We need to consider the following Cases (i) and (ii).
Case (i): $|T^{L}\cap[i_{1}+1,i_{d}]|=1$. Then by the conditions of Case 2, and by (13), we have $k_{1}\leq j\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}$ (see Fig. 4 (a)). Similar to Claim 1.1, we can prove $x_{[k_{1},i_{1}]}~{}($resp. $x^{\prime}_{[i_{d},k^{\prime}_{1}-1]})$ is contained in a run of $\bm{x}~{}($resp. $\bm{x}^{\prime})$, so we have $\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})=\{(x_{[n]\backslash j},x^{\prime}_{[n]\backslash j^{\prime}}):k_{1}\leq j\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}\}=\{(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash i_{d}})\}$.
Case (ii): $|T^{L}\cap[i_{1}+1,i_{d}]|=0$. In this case, by conditions of Case 2, and by (13), we have two possibilities: 1) $k_{b}\leq j<k_{1}\leq i_{1}<i_{d}\leq j^{\prime}<k^{\prime}_{1}$ (see Fig. 4 (b)), which implies $(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}})\in\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})$; and 2) $k_{1}\leq j\leq i_{1}<i_{d}<k^{\prime}_{1}\leq j^{\prime}\leq k^{\prime}_{b}$ (see Fig. 4 (c)), which implies $(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\in\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})$. Thus, we have $\Lambda^{L}_{1,2}(\bm{x},\bm{x}^{\prime})\subseteq\{(x_{[n]\backslash k_{1}-1},x^{\prime}_{[n]\backslash i_{d}}),(x_{[n]\backslash i_{1}},x^{\prime}_{[n]\backslash k^{\prime}_{1}})\}$.
Refer to captionRefer to caption
Figure 4: An illustration of the pair $(j,j^{\prime})$ in the proof of Claim 1.2. Here $S=\{i_{1},i_{2},i_{3}\}$. In this figure, (a) is for Case (i), (b) is for possibility 1) of Case (ii) and (c) is for possibility 2) of Case (ii). Here, $k_{1},k_{1}^{\prime},k_{b}$ and $k_{b}^{\prime}$ are defined by (7), (8), (11) and (12) respectively.
@@
Clearly, $(j,j^{\prime})$ satisfies the conditions of Case 3 ((((i.e., $|S\cap[j^{\prime}+1,n]|=1$ and $|S\cap[1,j-1]|=|T^{L}\cap[j+1,j^{\prime}]|=0)$ if and only if (see Fig. 5)
@@
Figure 5: An illustration of the pair $(j,j^{\prime})$ in the proof of Claim 1.3. Here, $S=\{i_{1},i_{2},i_{3}\}$ and $j_{1}^{\prime}=\min(T^{L}\cap[i_{2}+1,i_{3}])$ because $d=3$.